A roller coaster ride in a park includes a vertical drop of 60 m. Assume that the roller coaster has a speed of nearly zero when it reaches the top of the hill. Neglecting friction, find the speed of the riders at the bottom of the hill

1
Expert's answer
2021-11-18T15:09:23-0500

mgh=mv2/2v=2gh=29.860=34.3 (m/s)mgh=mv^2/2\to v=\sqrt{2gh}=\sqrt{2\cdot 9.8\cdot 60}=34.3\ (m/s) . Answer